OpenStudy (anonymous):

(16-9x^2) ^(3/2)/ x^6

OpenStudy (anonymous):

please help

OpenStudy (campbell_st):

my only though it in the numerator \[\frac{(16 -9x^2)^{3/2}}{x^6} = \frac{(4-3x)^{3/2}(4 + 3x)^{3/2}}{x^6}\]

OpenStudy (campbell_st):

or as an alternative \[\frac{\sqrt{(16 - 9x^2)^3}}{x^6}= \frac{(16-9x^2)\sqrt{(16-9x^2}}{x^6}\]

OpenStudy (anonymous):

whats next then?

OpenStudy (campbell_st):

ok... this can be a split fraction \[\frac{16\sqrt{16 - 9x^2}}{x^6} - \frac{9x^2\sqrt{16 - 9x^2}}{x^6}\] eliminate the common factor of x^2 in the 2nd term leaves \[\frac{16\sqrt{16 - 9x^2}}{x^6} - \frac{9\sqrt{16-9x^2}}{x^4}\] I think the original form is a lot friendlier

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+of+%2816-9x%5E2%29+%5E%283%2F2%29%2F+x%5E6 please do check this solution is this right?

OpenStudy (campbell_st):

lol... just trust wolframalpha

OpenStudy (anonymous):

ok